OpenStudy (anonymous):

How many solutions does the equation have? 3(4x-2)=2(6x-3) A. one solution B. infinite solutions C. no solution

OpenStudy (just_one_last_goodbye):

Give me a sec ok? ^_^

OpenStudy (anonymous):

okay

OpenStudy (just_one_last_goodbye):

You want to find out if the problem is true or false correct?

OpenStudy (just_one_last_goodbye):

There are 2 sides to the problem lets start with the left

OpenStudy (anonymous):

ok

OpenStudy (just_one_last_goodbye):

3(4x -2) lets remove parenthesis and multiply 3 * 4x -2 * 3

OpenStudy (just_one_last_goodbye):

Ok so whats 3 * 4x = ?

OpenStudy (anonymous):

12x? @just_one_last_goodbye

OpenStudy (just_one_last_goodbye):

correct ^_^

OpenStudy (just_one_last_goodbye):

2 * 3 = ?

OpenStudy (anonymous):

6

OpenStudy (just_one_last_goodbye):

Good so the equation now looks like this ^_^ 12x - 6 = 2(6x - 3)

OpenStudy (just_one_last_goodbye):

now lets do the right side

OpenStudy (anonymous):

ok

OpenStudy (just_one_last_goodbye):

2(6x - 3) so... 2 * 6x -2 * 3

OpenStudy (just_one_last_goodbye):

2 * 6x = ?

OpenStudy (anonymous):

12x

OpenStudy (just_one_last_goodbye):

Alright and 2 * 3?

OpenStudy (anonymous):

6

OpenStudy (just_one_last_goodbye):

Good so both sides of the equation are 12x - 6 = 12x - 6 ^_^ meaning they are both equal and their true

OpenStudy (just_one_last_goodbye):

Hope this helps you

OpenStudy (anonymous):

thank you!

OpenStudy (just_one_last_goodbye):

No prob Please medal me ^_^ if you find my steps helpful